• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name sarty

    Log entry time 00:05:45 on July 24, 2006

    Entry number 175079

    keyword=Shift summary - with Run info for Kin 2 (LK2 and Ta targets)

    16:00 Start shift when LD2 target is being put in place to start next production run. Notice that right arm Q1 power supply has tripped off with a new error message. Will call in Scott, but continue data-taking with production on left arm.

    17:00 request access to fix HRS-R Q1 power supply fault (and move to Carbon target) -> radiation monitor in BSY tunnel trips immediately after we get to Controlled Access, so RADCON has to come in and fix/replace that monitor during our Access.

    19:00 Back to Beam permit...plan to continue Kin 2 (Carbon). BUT beam never comes back. MCC reports vacuum problem somewhere in the machine, and by 23:00 they report that the experts to check out the problem will be in early Monday morning...so beam is off until sometime Monday.

    --------------------------------------- RUN INFO FOR Kin 2 Production:

    LD2 - Run 1601 (deadtime too high: singles rate ~ 1.2 kHz - JUNK Run 1602 (prescale=3, now deadtime=0...take data @ ~300 Hz DAQ) 132 k events Run 1603 (prescale=2, deadtime=1%..take data @ ~ 600 Hz DAQ) 88 k events

    stop to move targets (beam is up and down, so take advantage of down) moving target @16:45

    Ta - Run 1604 (rate only ~300 Hz, but had prescale=2) abort JUNK Run 1605 (prescale=1, DAQ rate ~ 380 Hz, deadtime ~ 5%) 134 k events

    17:00 request access to fix HRS-R Q1 power supply fault (and move to Carbon target)